Caroline is 12 years older than Nick. Mark is 4 years younger than Nick. If the total of their ages is 44, how old is the youngest of them?

Answer by Cromlix(4381)

Hi there,
Make Nick's age = x
Caroline's age = x + 12
Mark's age = x - 4
x + (x + 12) + (x - 4)
Remove brackets
x + x + 12 + x - 4 = 44
Collect like terms
3x = 44 - 12 + 4
3x = 36
x = 12
Mark is 8 yrs old
Nick is 12 yrs old
Caroline is 24 yrs old
Youngest is Mark = 8 years old.
Hope this helps :-)